OpenStudy (anonymous):

A parachutist's rate during a free fall reaches 198 kilometers per hour. What is this rate in meters per second? At this rate, how many meters will the parachutist fall during 10 seconds of free fall? Do not round your answers.

OpenStudy (anonymous):

55 meters per second. Use 1 kmph=0.27 mps Distance traveled in 10 seconds = 55 x 10 = 550 meters

OpenStudy (unklerhaukus):

\[198\,[\text{km/hr}]\\ =198\,[\text{km/hr}]\times1\times 1\\ =198\,[\text{km/hr}]\times\frac{[\text{hr}]}{3600\,[\text s]}\times\frac{1000\,[\text m]}{[\text{km}]}\\ =\]

OpenStudy (unklerhaukus):

\[=\left(198\times\frac1{3600}\times 1000\right)\,[\text{m/s}]\\ =\]

OpenStudy (unklerhaukus):

for the second part of the question, use \[v = d/t\] (solve for \(d\))